A projectile is launched from rest at x = 0 and moves along a parabolic path described by y = 0.01 x2 m. If the y component of acceleration is ay = 0.1 t 3 m/s2 , where t is in seconds, calculate the magnitude of the projectile’s velocity and acceleration when t = 5 s.
Answer #1

ay(t)=0.1t3 m/s2

at t=5s,

ay(t=5s)=0.1(5)3 m/s2= 12.5 m/s2

v(t) =v0+a(t)*t

Here, v0 is the velocity at initial point.

Since initially the projectile is at rest, v0=0m/s

At t=5s

v(t=5s) =v0+a(t=5s)*t

v(t=5s) =(0m/s)+ 12.5 m/s2*5s

v(t=5s) =62.5 m/s